the park is a perfect right triangle. its hypotenuse is 40m long one leg of the triangle is 8 m shorterthan the other leg. find the length of each leg of the triangle.
Answer by stanbon(75887)

the park is a perfect right triangle. its hypotenuse is 40m long one leg of the triangle is 8 m shorterthan the other leg. find the length of each leg of the triangle.
-------------
Let one of the legs be "x" ; Let the other leg be "x-8"
Use Pythagoras to get:
x^2 + (x-8)^2 = 40^2
x^2 + x^2-16x+64 = 1600
2x^2-16x+64 = 1600
x^2-8x-768=0
(x-32)(x+24)=0
Positive answer:
x= 32 meters (the longer leg)
x-8 = 24 meters (the shorter leg)
Cheers,
Stan H.